Suppose 𝑓 is odd and differentiable everywhere. Prove that for every positive



number 𝑏, there exists a number 𝑐 in (βˆ’π‘, 𝑏) such that 𝑓′(𝑐) = 𝑓(𝑏)/𝑏.

1
Expert's answer
2022-03-29T11:39:20-0400

ANSWER.

The function "f" is differentiable everywhere, hence "f" is continuous everywhere.Let "b>0" , so "f" is differentiable on a closed interval "[-b,b]" ( hence "f" is continuous on "[-b,b]" ). Then , by Lagrange's Mean Value Theorem , there exists a point "c\in(-b,b)" , such that

"f(b)-f(-b)=f'(c)\cdot[b-(-b)]=2b\cdot f'(c)" . (1)

Since "f" is odd, so "f(-b)=-f(b)" or "f(b)-f(-b)=2f(b)." From (1) we get

"2f(b)=2b\cdot f'(c)."

Therefore, for every "b>0" there exists "c\in(-b,b)" , such that

"f'(c)=\frac{f(b)}{b}" .